KSU Summer 2007


Thursday May 31st  began project – got a tour of the lab, began reading about MOCVD in the book “The Blue Laser diode”.


Friday – June 1st.  Read articles and more from the book. 


Monday June 4th.  I think this was when I got shown the MOCVD machine in more detail.  Also read


Tuesday – got to see PL machine in action, but not in detail. 


Wednesday – got to see X-ray machine in detail. 


Thursday – got to see the PL machine in detail.


Friday – saw AFM in detail.  Got suggestion to work on NSOM


Monday – Began reading about NSOM


Tuesday – turned on some machines, began to look at it.  Didn’t really do anything to NSOM


Wednesday – turned on all of the machines.  Didn’t actually move the head or anything, but checked that initial procedure had been done.  No real adjustments. 


Thursday – saw AFM in much more detail.  Could probably operate it now.


Friday – was going to use AFM but Chris got sick


Monday – Didn’t really accomplish that much.  Re-read NSOM manual


Tuesday – Did more reading.  Not much else. 


Wednesday – Used the AFM!  Chris was watching but I was doing.


Thursday – I was going to use the AFM on my own today, but I didn’t get samples.


Friday – AFM is broken.  It’s getting sent to the manufacturer for repairs. 


Monday – Categorized the contents Drawer #5 (NSOM equipment)


Tuesday – Finished Drawer #5, moved on to NSOM drawer (More NSOM equipment)  also played with the NSOM some more


Wednesday – I found the resonance frequency for our NSOM head.  It was very close to the predicted value


Thursday – Got ready to bring NSOM online.  Found the actual NSOM frequency (I was wrong yesterday).  Broke the NSOM head. 


Friday – Found a new NSOM head, and hooked it up.  Got NSOM into contact mode. 

              When we tried to run the linearization program, the program crashes.  AFM got back from repairs


Monday – Fixed crashing problem and ran linearization.  Then took sample thing to look at it under the AFM. 

      It looks really weird, so we must find a better place to look at.  Took pictures with the Optical microscope to compare. 


Tuesday – Took images with the AFM.  Got three different places to look at.


Wednesday – Made the pictures all easy to look at, and brought them to my computer.


Thursday – Began trying to take pictures with NSOM, but stuff doesn’t look right.  Also, NSOM is crashing for linearization again. 


Friday – Well, the NSOM  seems to not be doing anything right.  The features it sees are 20 times bigger than what the AFM saw, and has no detail. 

   Also, I saw that it scratching the surface, and when I went to check it under the AFM I saw that I had dug a big trench in the metal sample. 

   When I started looking at the cables, I found one that was broken, but I don’t think it could cause all of the problems we saw. 


Monday – Found the part we needed, but the company asked me to email them, and I haven’t heard back yet.


Tuesday – Chris is trying to order the part.  I looked at the NSOM more, and I think the one wire could be causing (in theory) all of the problems. 

                  What I thought was the problem wasn’t.  Took optical microscope pictures of the trenches the NSOM dug.  Ordered the parts.


Wednesday – The parts shipped today.  Should be in Friday.  I took a look at the pictures Jagat had taken on the NSOM, and analyzed them. 


Thursday – Learned about Fourier transforms.  Actually came in late today. 


Friday – Parts didn’t come in.  Chris was apparently in the basement today.  Worked on the webpage.


Monday – Part came.  Got wire fixed.  It didn’t do anything different.  Crud.  Chris called Nanonics, they told us to email them, so I did. 

                  Didn’t get an email back today.  Tried lots of stuff on the NSOM but nothing helped. 


Tuesday – Got the email from Nanonics.  I sent an email back.  Tried a bunch of stuff that they suggested. 


Wednesday – Tried more stuff on the NSOM.  At the end of the day, the approach/retract has stopped working.


Thursday – Got the approach retract working, took three pictures, and now the approach retract has stopped again.


Friday – Tried unsuccessfully to get the approach/retract working, but to no avail.  Email Nanonics. 


Monday – Got an email from Nanonics and emailed them back.  Chris and I fixed the approach/retract ourselves, with help from the machine shop.

                  Didn’t have time to really take any pictures. 


Tuesday – Took a few pictures.  They looked a lot better, but took a lot longer to take.  The approach/retract failed at the end of the day.


Wednesday – Tried to fix the approach/retract today, but we were unsuccessful.


Thursday – Got the approach/retract fixed by the machine shop.  Took AFM picture of a new sample.  Took AFM pictures with the NSOM. 

                    The images look better, but not great. 


Friday – Took a whole bunch of pictures.  Some looked really nice, but they take a long time to take. 

   Still not sure what a Photonic crystal would look like, though. 


Monday – Took pictures of a similar structure to a photonic crystal.  It looks good.  We just hope the NSOM works.


Tuesday – Got a sample to look at.  It has periodic holes in it from growth errors.  We want to see the optical properties of these holes. 


Wednesday – Took NSOM images!  They look very different from the surface, indicating effects from underneath.  Wasn’t quite what we expected.

                        All the NSOM images look pretty much the same.  But it does seem to be working.